OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 5 and sqrt2 2,sqrt3 3sqrt3 sqrt29

OpenStudy (anonymous):

so you know the Pythagorean theorem is: a^2+b^2=c^2

OpenStudy (anonymous):

5 root 2 is a and b

OpenStudy (anonymous):

i got 3sqrt 3 am i right ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think 5 root 2 squared is 50. so 50 + 50=100

OpenStudy (anonymous):

im doing something wrong

OpenStudy (anonymous):

@camerondoherty @ShortyB can u guys help. I can't even do this simple problem

OpenStudy (anonymous):

ok @lacrosseplayer22 is right you do have to use the Pythagorean theorem

OpenStudy (anonymous):

5 root 2 squared is 50 right?

OpenStudy (anonymous):

but 5^2 is 25 because 5 x 5 is 25 not 50

OpenStudy (anonymous):

no 5 ROOT 2 not 5 squared

OpenStudy (anonymous):

\[5\sqrt{2}\]

OpenStudy (anonymous):

none of my choices though

OpenStudy (anonymous):